### v1

# Grounded 3D-LLM with Referent Tokens

Abstract
Prior studies on 3D scene understanding have primarily developed specialized models for specific tasks or required task-specific fine-tuning. In this study, we propose Grounded 3D-LLM, which explores the potential of 3D large multi-modal models (3D LMMs) to consolidate various 3D vision tasks within a unified generative framework. The model uses scene referent tokens as special noun phrases to reference 3D scenes, enabling the handling of sequences that interleave 3D and textual data. It offers a natural approach for translating 3D vision tasks into language formats using task-specific instruction templates. To facilitate the use of referent tokens in subsequent language modeling, we have curated large-scale grounded language datasets that offer finer scene-text correspondence at the phrase level by bootstrapping existing object labels. Subsequently, we introduced Contrastive LAnguage-Scene Pre-training (CLASP) to effectively leverage this data, thereby integrating 3D vision with language models. Our comprehensive evaluation covers open-ended tasks like dense captioning and 3D QA, alongside close-ended tasks such as object detection and language grounding. Experiments across multiple 3D benchmarks reveal the leading performance and the broad applicability of Grounded 3D-LLM. Code and datasets will be released on the project page.
1 Introduction
The pursuit of a unified model for 3D scene understanding, capable of performing various vision tasks, is a longstanding and significant challenge. Existing 3D scene-level models are typically customized for specific downstream tasks, including 3D object detection [37, 50, 65, 62], language grounding [8, 84, 34, 9, 12], question answering [4, 20, 52, 39], and dense captioning [14, 15, 67, 9, 6, 13, 39]. Recent 3D large multi-modal models (LMM), such as Chat-3D [72] and LL3DA [11], facilitate visual interactions based on detected objects for text reasoning, yet their capabilities remain confined to text-level tasks. 3D-LLM [31], which extends 2D VLMs [43] to 3D, grapples with challenges in 3D spatial reasoning due to the intrinsic knowledge derived from perspective views. This results in suboptimal performance in crucial localization tasks like object localization and language grounding, thereby limiting its practical applicability. The capacity of large multi-modal models (LMMs) to pinpoint objects or regions in complex environments is vital, especially for applications such as VR/AR, robotics, interactive embodied agents, object navigation, and manipulation.
However, the exploration of a unified model for 3D perception and reasoning remains scarce, leaving it as an open problem. In our study, by leveraging the generative modeling capability of large language models (LLM), we introduce Grounded 3D-LLM for integration of multiple 3D vision tasks in linguistic formats (Fig. 1). To connect perception and language, this model represents scene regions or object features as a special token, termed “referent tokens” (<ref>). Referent tokens support decoding to identify their corresponding entities in the scene. For instance, <ref> can denote entities such as “the nearby nightstand” or “three brown chairs”, substituting the corresponding text tokens in the language. The use of referent tokens as manageable noun phrases in linguistic descriptions includes two main features: (i) A referent token can either replace or accompany text phrases, as depicted in Fig. 1. (ii) A referent token can be decoded into scene entities, encompassing single or multiple objects in a 3D scene, using simple decoder heads. With this design, referent tokens serve as the unified interface, facilitating 3D grounding capabilities in various 3D vision tasks.
The referent token in our language modeling relies on phrase-level scene-text correspondence, setting it apart from models like CLIP [57], which align at the sentence level. To achieve this, inspired by phrase-to-region correspondence [40], we employ Contrastive LAnguage-Scene Pre-training (CLASP) for a more granular scene-text correspondence. Specifically, CLASP aligns noun phrases with visual queries [62] representing scene objects or regions. However, large-scale paired scene-text data at the phrase level remains scarce. To address this, we developed an automated pipeline for annotating scene-level language data, utilizing densely-annotated object labels from indoor detection datasets to create the Grounded Scene Caption (G-SceneCap) data. This dataset comprises 107K local scene captions with grounding, offering high object density with an average of one instance every nine tokens as shown in Table 2. Additionally, we converted existing human-annotated language data [8, 83] into grounded language forms for training and evaluation. Our experiments using this data in supervised and zero-shot settings [88] show broad generalization across grounding benchmarks.
After the pre-training stage, we conduct multi-task instruction tuning for Grounded 3D-LLM using referent tokens. This model supports the decoding of these tokens into mask embeddings, enabling the retrieval of point-level masks and thus identifying their external scene referents. It supports a range of existing 3D vision tasks—from single to multi-object grounding, instance segmentation, to 3D QA and captioning—by employing diverse templates of task-specific instruction-tuning, as shown in Fig. 1. We compare this model’s prompt support and task versatility with previous large multi-modal models in Table 1. We believe this represents a valid step toward unifying various 3D vision tasks within generative language modeling.
Figure 1: Unifying scene-level 3D vision tasks via task-specific prompts in Grounded 3D-LLM framework. The interleaved referent token <ref> in the language output enables decoding for point-level masks. Matching colors are shown in the figure to connect phrases with their respective object masks.
To summarize, our contributions are as follows:
  • We developed Grounded 3D-LLM, which first establishes a correspondence between 3D scenes and language phrases through referent tokens. This method enhances scene referencing and effectively supports 3D vision tasks in language modeling, including single- and multi-object grounding, along with introducing 3D detection for the first time.
  • We developed a carefully crafted automated 3D scene caption dataset curation pipeline that provides finer correspondence at the phrase level. Experiments using CLASP in both supervised and zero-shot text settings demonstrate the effectiveness of pre-training on this data for phrase-level scene-text alignment.
  • The Grounded 3D-LLM model tackles 3D grounding and language tasks generatively without the need for specialized models. It achieves top-tier performance in most downstream tasks among generative models, particularly in grounding problems, without task-specific fine-tuning.
Method LLM Prompts Tasks
Text Vision Inst.Seg. Obj.Det. Grd. Point-Grd. Multi-Obj Grd. QA Cap
PointGroup [37]
Mask3D [62]
Multi3DRef [83]
BUTD-DETR [36]
3D-VisTA [88]
Chat-3D [72]
Chat-3D v2 [33]
3D-LLM [31]
LL3DA [11]
Grounded 3D-LLM
Table 1: Comparison of 3D multi-modal models. We refer to instance segmentation as Inst.Seg., object box detection as Obj.Det., single-object grounding as Grd., point-level grounding as Point-Grd., multi-object grounding as Multi-Obj Grd., question answering as QA, and dense captioning as Cap.
2 Related Works
2D large multi-modal models. Vision-language models [57, 44, 43] are proposed to bridge visual and textual models via image-level contrastive learning. Their direct application enables open-world understanding [28, 80]. GLIP and its subsequent works [45, 82, 77, 49] have innovatively formulated object detection as grounding problems, leveraging detection or pseudo-labels to align semantics. With stronger language models, various powerful large multi-modal models are capable of handling tasks including image captioning [35, 54, 3], visual question answering [47, 87, 17, 70], and object detection/segmentation [54, 49, 58]. Besides, incorporating multi-modal tokens into language modeling has been validated, including visual understanding [3, 69, 42, 81], robot control [59, 5], and interleaved text-image content generation [63, 25, 26].
3D scene understanding models. In the domain of 3D scene comprehension, language is progressively utilized as a contextual tool to encapsulate user intentions, given its extensive applicability. This spans from close-set object detection [37, 50, 62, 68, 36] to open-set 3D detection [10, 22, 23, 86]. Typically, these 3D detectors lay the groundwork for subsequent scene-language tasks. Such tasks encompass 3D Visual Grounding [8, 2, 34, 83, 12, 71, 84], where models pinpoint objects within a 3D scene based on language queries. They also include 3D Question Answering [4, 53, 51], which demands robust spatial perception and reasoning. Another task, 3D Dense Captioning [14, 79, 38], involves localizing and describing all scene objects, necessitating a comprehensive understanding of object locations and attributes. While initial efforts were often task-specific and lacked generalizability, some methodologies [6, 9] have integrated 3D visual grounding and dense captioning tasks, leveraging their complementary nature. Recent endeavors such as 3D-VisTA [88] and 3D-VLP [39] aim to establish a universal 3D visual-language framework by pre-aligning 3D scenes with language. In contrast to our model, most still concentrate on close-set scene understanding, striving to develop specialist models or requiring per-task fine-tuning. Moreover, they typically rely on prior detectors.
3D large multi-modal models. 3D object-level LMMs [78, 74, 75, 48, 85] have effectively bridged the gap between 3D visuals and texts by leveraging extensive 3D object datasets [19, 18]. This has advanced the development of 3D object-level LMM [30, 29, 73, 31, 56, 55, 85, 46]. However, these models struggle with interpreting complex spatial interrelationships in 3D scenes. For scene-level LLMs, Chat-3D [72] and LL3DA [11] facilitate interaction with scene objects via pre-selection methods. Chat-3D v2 [33] enhances this by using object identifiers for precise referencing and grounding. 3D-LLM [31] improves scene understanding by incorporating positional embeddings and location tokens. Our approach utilizes special referent tokens for flexible 3D scene modeling and supports various tasks including object detection and grounding. Additionally, our model functions as a generalist model without requiring task-specific fine-tuning.
3 Method
This section is structured into three main parts. Initially, Sec. 3.1 explores different levels of vision-language alignment across sentences, words, and phrases. Next, Sec. 3.2 details the Grounded 3D-LLM, from vision-language pre-training to multi-task instruction-tuning using referent tokens. Finally, Sec. 3.3 describes construction of the grounded language data.
3.1 Preliminary: Vision-Language Correspondence
The significance of vision-language alignment in 2D foundation models, exemplified by the Contrastive Language-Image Pre-training (CLIP) model [61], is well-established. We explore three levels of vision-language alignment:
  • Sentence-level Correspondence: As exemplified by CLIP models [57, 44, 43] and 3D annotation works [8, 2, 83], this constraint aligns image-text pairs at the sentence level. However, it is limited by the scarcity of vision-language data, particularly in 3D.
  • Word-level Correspondence: This term describes the method of concatenating detection labels into a sentence, rather than embedding phrases within sentences, first introduced by GLIPs [45, 82, 77]. This approach reinterprets object detection labels as phrase grounding tasks. However, it typically lacks detailed appearance annotations and does not capture spatial relationships.
  • Phrase-level Correspondence: As referred to as phrase-to-region correspondence in prior work [40], this constraint requires linking text phrases to real-world objects within a context description. Correspondences may be one-to-one or multiple.
Our motivation. Phrase-to-region correspondence provides a more granular scene-text alignment than sentence-level correspondence, enabling effective modeling of grouping objects with similar attributes such as spatial relationships or appearances within the text context. A vision-language model is first pre-trained on extensive scene-text data to establish this phrase-level cross-modal correspondence. Utilizing this semantically enriched scene representation, an LLM model is then trained to integrate referent tokens into object clusters with shared linguistic descriptions. This process aligns the hidden embeddings of referent tokens with their corresponding visual referents, mirroring the phrase-to-region approach. For example, in phrases like “two brown chairs nearby <ref>” or simply “object <ref>”, a mask decoder can be trained to produce mask embeddings of <ref> to deduce the positions and shapes of these target chairs.
3.2 Grounded 3D-LLM
The main pipeline of Grounded 3D-LLM is depicted in Fig. 2. In the pre-training stage, contrastive language-scene pre-training (CLASP) is performed to align point and text embeddings at the phrase level on the extensive scene-text data. During the instruction-tuning phase, the overall Grounded 3D-LLM is trained to utilize referent tokens from the scene as soft text tokens to tackle various 3D vision tasks in a unified generative modeling framework.
Step 1: Contrastive Language-Scene Pre-training. Grounded 3D-LLM relies on referent tokens to reference the 3D scene, necessitating phrase-level alignment between natural phrases and scene referents. To bridge 3D and text modality gap, we propose the Contrastive Language-Scene Pre-training (CLASP). This model co-trains a 3D encoder, a text encoder, and a cross-modal interactor for text-guided instance segmentation as the left side of Fig. 2. In the CLASP’s implementation, a sparse convolutional network [27] serves as the 3D encoder to extract point-level embeddings $P$, while BERT [21] is utilized for extracting text embeddings $T$. Utilizing these embeddings, the cross-modal interactor employs $K$ learnable queries $Q$ as proxies to connect text with raw point clouds. These queries interact with point embeddings through cross-attention modules and with text embeddings via bidirectional cross-attention modules [49]. The model is initialized with a pre-trained 3D detector [62] to accelerate the convergence of queries serving as object proposals. At last, this multi-modal model outputs two text embeddings: $P_{text}$ for positive text tokens and $Q_{text}$ for scene queries, along with $P_{mask}$ and $Q_{mask}$, the mask embeddings for points and queries, respectively:
$\left[P_{mask},Q_{mask}\right],\left[Q_{text},P_{text}\right]=\text{Cross-Modal Interactor}(P,T,Q).$ (1)
As depicted in Fig. 2, we compute the alignment scores $S_{mask}$ for per-point mask classification and similarity $S_{text}$ between text phrases and their corresponding entities (could be multiple instances):
$S_{mask}=P_{mask}Q_{mask}^{\top},\ \ S_{text}=P_{text}Q_{text}^{\top}.$ (2)
The contrastive loss for $S_{mask}$ is calculated based on assignments using bipartite matching. Upon establishing the matching indices for the mapping between instances and queries, another contrastive loss for noun phrases and instance alignment, $S_{text}$ is supervised. Both similarity matrices provide mappings from text phrases to point-level masks.
Figure 2: Main pipeline for Grounded 3D-LLM. The training process for Grounded 3D-LLM is divided into two key steps. Firstly, CLASP utilizes extensive scene-text annotations (at the phrase level) to pre-train a 3D point cloud encoder and a cross-modal interactor. The subsequent step involves multi-task instruction tuning, which interlaces referent tokens within the instructions and responses, thereby facilitating flexible 3D scene understanding tasks.
Step 2: Instruction tuning with scene referent tokens. During instruction tuning, the model learns to use referent token as the noun phrase to reference the scene, represented by a special token <ref> in the language. The language input and output formats are demonstrated in the following example:
[All Scene queries] USER: Describe the object <ref> in the scene.
ASSISTANT: <p> A black chair </p> <ref> with four legs.
In this format, [All scene queries] derive from projecting Q-Former’s query features into the language embedding space. For input, <ref> is filled with the soft query embeddings via instance-query mapping. For output, <ref> serves as a placeholder token, aligning its hidden embeddings $R_{text}$ with the corresponding scene entities (single or multi-instances) as calculated by $S_{text}$:
$S_{ref}=R_{text}Q_{text}^{\top}.$ (3)
During training, the contrastive loss for $S_{ref}$ is applied as $S_{text}$ in the pre-training stage. During inference, referent tokens are decoded with a small decoder (two MLP layers) that extracts mask embeddings, retrieves queries and obtains their associated masks. Through the above definition of <ref> in language inputs and outputs, it enables a closed-loop dialogue with grounding in the 3D scene. To enhance the readability of the referent tokens, the model is trained to generate relevant text phrases followed by the tokens, such as “<p> three nearby chairs </p> <ref>.” special token pairs <p> and </p> signify the start and end positions of the noun phrases [58, 81]. If no category is provided, we use the generic term “object” instead.
Task-specific instruction templates. The flexible design of referent tokens enables the unification of downstream 3D vision tasks into language forms. We transform existing 3D vision tasks into instruction formats, as depicted in Fig. 1. This format encompasses single and multi-object grounding [8, 83], object detection [16, 60], dense captioning [14], 3D QA [4], and descriptions of objects and local scenes. For each task, we utilize a range of task-specific instruction templates (about 10-20), tailored to the data annotation formats. For tasks requiring grounding or detection, we append “(with grounding)” to instructions to indicate the need for generating referents in language outputs. These templates can be found in the supplementary materials.
Discussion: Referent token as localization token. Prior work 3D-LLM [31] employs discrete location tokens, which constrain the granularity of localization accuracy. Moreover, it lacks point-level grounding capabilities and struggles with grounded visual chat tasks. Chat-3D v2 [33], pairing object identifiers and object features as text inputs, confronts high token costs and dependency on independent detectors. Related works in 2D LLM [42, 81] rely on downstream expert models [41] to improve localization. Differently, our model operates without a separate detector and object feature extractor, relying instead on CLASP that aligns semantics with scene queries. It enables phrase-to-region mapping and associated mask retrieval. Additionally, each referent token can be decoded into multiple instances, thereby reducing token costs.
3.3 Grounded Language Data
In this work, we define “grounded language” as language annotations that exhibit phrase-to-region correspondence, as detailed in Sec. 3.1. We adopt two approaches to generate grounded language data: (i) automatic generation of grounded scene captions, and (ii) transformation of existing language data into a grounded format. Compared to existing datasets [8, 83, 2], the automated method offers two advantages: First, the captioning employs all objects as anchors to ensure complete coverage of 3D scenes. Second, the generated caption has a high object density, averaging one instance per nine tokens, as shown in Table 2, ensuring the training efficiency of positive instances per text.
Figure 3: Grounded scene caption dataset generation pipeline. Matching colors are used to link phrases with their corresponding object masks.
Grounded scene caption data generation. The dataset generation for the Grounded Scene Caption dataset (G-SceneCap) leverages ChatGPT and 2D vision-language models, utilizing dense object annotations from existing 3D scan datasets [16, 66, 7], as illustrated in Fig. 3:
Step 1: Bootstrapping object captions with GT label correction. Using 3D real-scan datasets, we annotate each object with the vision-language model CogVLM [70], using the images of the largest visible areas. Inconsistent annotations are rectified using raw instance labels.
Step 2: Condensing objects in local scenes into a caption. For each enumerated anchor object, we form an initial object set by randomly selecting a group of nearby objects. Their captions and coordinates $(x,y,z)$ are input into GPT-4 for captioning, which requires referencing objects by their IDs in the format “[object_phrase object_ID]” in the caption.
Step 3: Adding rule-based relations. To enrich scene captions, we integrate program-generated spatial relationships from Sr3D [2]. By selecting an anchor object from the set in step 2, we apply the spatial relation rules (e.g., between, supporting, nearest, back) to include related objects. GPT-4 then combines these relationships into the prior caption from step 2.
Conversion of existing data. In this work, we employ ChatGPT [1] to extract positive phrases for ScanRefer and Multi3DRef datasets for training and evaluation. We refer to the phrase-grounding datasets as “Grounded ScanRefer” (G-ScanRefer) and “Grounded Multi3DRefer” (G-Multi3DRefer). Additionally, we define “Grounded Detection” (G-Detection) as the transformation of detection labels into a sequence of text prompts, similar to GLIP [45].
Extension to embodied dialogue and planning. Grounded scene captions provide local scene information. Using these, we prompt GPT-4 to generate multi-round dialogues and detailed planning sequences in grounded language format. These dialogues, between a human and an embodied agent, encompass tasks including completing activities, discussing layouts, and conducting object-oriented QA. The data generation prompts are detailed in supplementary files.
Table 2: 3D scene-language dataset comparison. We calculate the ratio of instances to tokens to assess training effectiveness per sentence, using the BERT tokenizer [21] to determine the average token length. For sentences with single-object correspondence, each text is paired with one instance.
Dataset #3D Scan #Language Correspondence Average Tokens Inst/Token
Text-level Object-level
ScanRefer [8] 0.7K 52K Sentence Single 22.6 4.4%
Sr3D [2] 1.5K 84K Sentence Single 13.1 7.7%
Sr3D+ [2] 1.5K 115K Sentence Single 12.9 7.7%
Nr3D [2] 0.7K 42K Sentence Single 14.7 6.8%
Multi3DRefer [83] 0.7K 62K Sentence Multi 19.4 7.2%
ScanScribe [88] 3.0K 278K Sentence Single 66.5 1.5%
Grounded Scene Caption 1.5K 107K Phrase Multi 85.9 11.1%
4 Experiments
4.1 Experimental Setup
Datasets and benchmarks. Our evaluation is conducted on the ScanNet dataset [16], an extensive indoor 3D scene dataset that contains 1,513 multi-modal scenes with 3D point clouds, image sequences, and rich point-level instance segmentation annotations. The dataset is divided into 1,201 scenes for training and 312 scenes for validation, with all subsequent benchmarks adhering to these training/validation splits. The CLASP pre-training involves all grounded language data (Sec. 3.3), including G-SceneCap, G-ScanRefer, G-Multi3DRef, and G-Detection. We conduct our evaluation across several benchmarks: ScanRefer [8] and Multi3DRef [83] for single- and multi-object grounding, ScanNet-200 [60] for object detection, and ScanQA [4] and Scan2Cap [14] for 3D QA and dense captioning tasks. These datasets are converted to instruction formats for instruction-tuning and LLM evaluation. Detailed evaluation metrics are provided in the supplementary files.
Baseline methods. Existing methods for downstream tasks are split into discriminative and generative methods in Table 3. (i) Discriminative models perform closed-set classification via detected object proposals, covering single-object [12, 34, 84] and multi-object grounding [83]. 3D-VisTA [88] pre-trains a unified 3D backbone and conducts task-specific fine-tuning. The top 3D close-set detector, Mask3D [62] is shown for reference. (ii) Generative models cover recent 3D LMMs [72, 33, 11, 31], mainly necessitating task-specific fine-tuning for language grounding, termed Specialist LLMs. LL3DA [11] acts as the generalist, supporting language tasks but limited in localization.
4.2 Main Results
Table 3: Evaluation of 3D scene-level LMMs as a generalist. Entries in gray denote models specialized for specific datasets. The term “Generative Model” denotes generative models with open-ended outputs, unlike discriminative models that focus on closed-set classification. “Zero-shot” denotes models used directly without fine-tuning. “Specialist” and “Generalist” categorize models fine-tuned for particular tasks versus those trained jointly.
Method Generative Model ScanRefer Grd. Multi3DRef Grd. Scan2Cap ScanQA ScanNet-200 Det.
Acc@0.25
Acc@0.5
F1@0.25
F1@0.5
B-4@0.5
C@0.5
B-4 C AP AP0.25 AP0.5
ScanRefer [8] 37.3 24.3
MVT [34] 40.8 33.3
3DVG-Trans [84] 45.9 34.5
ViL3DRel [12] 47.9 37.7
M3DRef-CLIP [83] 51.9 44.7 42.8 38.4
Scan2Cap [14] 22.4 35.2
ScanQA [4] 10.1 64.9
3D-VisTA [88] 50.6 45.8 34.0 66.9 13.1 72.9
Mask3D [8] 27.4 37.0 42.3
Zero-shot LLMs
LLM-Grounder [76] 17.1 5.3
Specialist LLMs
3D-LLM(Flamingo)
[31]
21.2 7.2 59.2
3D-LLM(BLIP2-flant5)
[31]
30.3 12.0 69.4
Chat-3D [72] 6.4 53.2
Chat-3D v2 [33] 35.9 30.4 7.3 77.1
LL3DA [11] 36.8 65.2 13.5 76.8
Generalist LLMs
LL3DA [11] 36.0 62.9 13.3 75.7
Grounded 3D-LLM 47.9 44.1 45.2 40.6 35.5 70.6 13.4 72.7 15.1 20.8 23.5
Table 4: Comparison of CLASP with previous referring and detection methods. “Point-supervision” involves using point-level instance segmentation supervision. “End-to-end models” are those that can be trained either fully end-to-end or using separate detectors.
Method Point- supervision End-to -end ScanRefer Grd. Multi3DRef Grd. ScanNet200 Det.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 AP AP@0.25 AP@0.5
BUTD-DETR [36] 52.2 39.8
ViL3DRel [12] 47.9 37.7
3D-VisTA [88] 50.6 45.8
M3DRef-CLIP [83] 51.9 44.7 42.8 38.4
Mask3D [62] 27.4 37.0 42.3
CLASP 53.2 46.7 51.5 47.3 27.2 36.4 42.1
2×2$\times$2 × schedule 54.3 47.4 54.6 49.6 27.5 37.0 41.4
Grounded 3D-LLM. As shown in Table 3, Grounded 3D-LLM, without task-specific fine-tuning, outperforms previous LLM-based models in most metrics, positioning it as a promising candidate for a unified framework in 3D scene understanding. Compared to previous generative models, without task-specific fine-tuning, it achieves an improvement of over 10 points in both single and multiple object grounding. Additionally, it attains comparable outcomes in language-based tasks, achieving the highest CIDEr score of 70.6 in Scan2Cap and matching the top BLEU-4 score of 13.4 in ScanQA, surpassing previous method LL3DA, which is limited to language output without referencing capabilities. Notably, it is the first 3D generative model capable of 3D object detection, distinguishing it from its predecessors.
CLASP. As shown in Table 4, CLASP excels in 3D grounding and detection benchmarks, especially in ScanRefer grounding, Multi3DRef multi-object grounding, and ScanNet-200 detection, highlighting its enhanced phrase grounding capabilities compared to earlier models. Notably, in the Multi3DRef grounding task, CLASP outperforms specialized models by over 8 points, demonstrating its strong multi-object handling. With a doubled training schedule, CLASP exceeds all rivals in grounding accuracy and matches the Mask3D model in detection performance. Note that in this study, we primarily use the 1×1$\times$1 × model for Grounded 3D-LLM, as the 2×2$\times$2 × model, while better at grounding, falls short in subsequent language tasks.
4.3 Ablation Studies
Table 5: Ablation studies on the grounding ability of Grounded 3D-LLM. The term “all data” encompasses the G-ScanRefer, G-Multi3DRef, and G-SceneCap datasets used for pre-training.
# CLASP Pre-training data Output Referent type ScanRefer Grd. Multi3DRef Grd.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5
(a) $1$-to-Many 6.5 5.8 6.4 5.3
(b) G-ScanRefer $1$-to-Many 43.2 36.6 40.2 36.9
(c) G-SceneCap $1$-to-Many 47.6 43.4 44.3 40.1
(d) All data $1$-to-$1$ 46.5 43.3 44.2 40.7
(e) All data $1$-to-Many 47.9 44.1 45.2 40.6
Table 6: Effects of contrastive language scene pre-training for phrase grounding. $\dagger$ represents using ground-truth boxes.
# Pre-training Method Correspondence Level Training Data ScanRefer Grd. Multi3DRef Grd.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5
(a) M3DRef-CLIP Sentence Multi3DRef 55.0
(b) 3D-VisTA Sentence ScanScribe 41.1 36.4
(c) (b) w. ScanRefer 50.6 45.8
(d) CLASP Phrase G-SceneCap 46.2 35.9 39.3 36.1
(e) (d) w. G-ScanRefer 49.7 44.4 39.2 35.9
(f) (e) w. G-Detection 51.1 44.8 40.2 37.0
(g) (f) w. G-Multi3DRef 53.2 46.7 51.5 47.3
(h) (g) w/o. G-SceneCap 51.1 45.4 48.5 43.1
Ablation studies on referent token. As shown in Table 5, comparing line (a) with others reveals that directly adapting existing closed-set detectors is ineffective for integrating referent tokens in LLMs. In contrast, CLASP pre-training enhances the interpretation of scene referents in language models. Additionally, we analyze two types of referent tokens: one-to-one (d), where each object is matched with a unique token through Hungarian matching, and one-to-many (e), where a single token represents multiple objects. The latter approach generally yields better results.
Dataset ablations for CLASP. Table 6 directly shows the phrase grounding capabilities of the CLASP model using different data across two key points: (i) Zero-shot text evaluation. Without training on ScanRefer or Multi3DRef language data, our CLASP model (d) proves effective compared to previous models (a) and (b). Notably, despite 3D-VisTA’s data generation adhering to sentence-level grounding similar to ScanRefer, our model exceeds it by 4.9 points in Acc@0.25. Furthermore, in experiment (d), our direct evaluation on the Multi3DRef dataset demonstrates robust multi-object grounding ability. (ii) Multi-dataset evaluation. We further investigate the effects of joint training with grounded language datasets, progressively integrating G-ScanRefer (e), G-Detection (f), and G-Multi3DRef (g). This integration boosts performance, achieving 53.2 in Acc@0.25 on ScanRefer and 51.5 in F1@0.25 on Multi3DRef. Notably, Comparison between (g) and (h) shows that the G-SceneCap dataset consistently raises performance by +2.1 points in ScanRefer Acc@0.25 and +3.0 points in Multi3DRef F1@0.25, highlighting its effectiveness.
5 Limitations and Conclusion
We explore the potential of large multi-modal 3D models to unify various downstream 3D vision tasks in an open-ended generative manner. By interpreting scene referents as language tokens, Grounded 3D-LLM links scene objects or regions with descriptions. We propose a contrastive scene-text alignment method that uses curated datasets to facilitate special token modeling in language. However, it exhibits limitations in performance compared to the pre-trained CLASP model, indicating a disparity between discriminative and generative approaches. Scaling the scene-text data to bridge this gap is an objective for future work.
Society Impact. The method enhances the grounding capabilities of large multi-modal models in 3D environments, generally benefiting downstream applications in robotics. However, it does not comprehensively address all indoor scenarios and may produce hallucinated language outputs, thereby posing risks to system integrity.
Supplementary Files for Grounded 3D-LLM
The supplementary materials for CLASP are organized as follows:
  1. 1.
    A detailed introduction to datasets and evaluation metrics is provided in Sec. A.
  2. 2.
    Details on pre-training and instruction-tuning implementations are outlined in Sec. B.
  3. 3.
    Additional ablation studies are included in Sec. C.
  4. 4.
    Visualizations of results and failure cases for downstream tasks are presented in Section D.
  5. 5.
    Details on grounded language data generation, along with visualizations of dataset statistics and spatial relationships, are illustrated in Sec. E.
  6. 6.
    Data generation prompts for the Grounded Scene Caption dataset are discussed in Sec. F.
  7. 7.
    Multi-task instruction templates are described in Sec. G.
  8. 8.
    Extension to embodied dialogue and planning data is shown in Sec. H
Appendix A Datasets and Evaluation Details
ScanRefer [8]: This benchmark evaluates the ability to localize 3D objects using natural language descriptions from the ScanNet dataset. We assess prediction accuracy by calculating the percentage of predictions with an Intersection over Union (IoU) above $0.25$ and $0.5$ compared to ground truth boxes, following the official metrics [8]. For a fair comparison with prior works, Grounded 3D-LLM extracts the object with the highest score according to the similarity scores $S_{ref}$, grounding a single object following previous studies [8, 88].
Multi3DRef [83]: This dataset is proposed to address the challenge of localizing a flexible number of objects in real-world 3D scenes using natural language descriptions on the ScanNet dataset. The evaluation for the Multi3DRef dataset [83] involves computing the F1 score at IoU thresholds of $0.25$ and $0.5$ between multiple predictions and ground-truth objects. The official metric uses the Hungarian algorithm to match predicted optimally and ground truth bounding boxes, based on the negative IoU as the cost function. For evaluation of Multi3DRef, we employ a score threshold of $0.3$ to filter the candidate objects.
ScanNet-200 Detection [60]: This dataset offers a comprehensive evaluation for 3D instance segmentation across $200$ classes. Evaluation metrics include mean average precision (AP) across different overlap thresholds: AP at $25\%$ overlap (AP@0.250.25{0.25}0.25), AP at $50\%$ overlap (AP@$0.5$), and AP averaged over overlaps in the range of $[0.5\text{:}0.95\text{:}0.05]$ (AP). These metrics follow the official ScanNet benchmark guidelines [16].
Scan2Cap [14]: The dense captioning benchmark is based on ScanRefer dataset [8], requiring models jointly perform detection and generate captions for each object. The evaluation of the language generation component includes BLEU-4 and CIDEr metrics, weighted by the Intersection over Union (IoU) score with the ground truth boxes, which must be greater than $0.25$ or $0.5$. Recent works [11, 6, 14, 13] employ different strategies for matching detection results with ground-truth annotations. We compute these language metrics following prior works [11, 13].
ScanQA [4]: This dataset is annotated based on ScanNet dataset for 3D Question Answering (3D-QA) task to evaluate the visual and spatial understand of 3D environments. This dataset facilitates the processing of visual data from 3D scans, enabling models to respond to text-based inquiries regarding these scenes. For evaluation, metrics of BLEU-4 and CIDEr metrics are utilized.
In the ScanRefer, Multi3DRef, and Scan2Cap benchmarks, bounding boxes within the axis-aligned coordinate system are typically utilized to calculate the 3D IoU [24]. For comparison with previous studies, we derive 3D bounding boxes by extracting the minimum and maximum values from the predicted masks. This post-processing is adopted even though our model provides a more precise point-level mask.
Appendix B Implementation Details
The training hyper-parameters are detailed in Table A-7. All models are trained on 8 NVIDIA A100 (80G) GPUs.
Table 7: Training hyper-parameters. The instruction fine-tuning batch size comprises 5 scenes, each containing up to 200 instructions.
Configuration CLASP pre-training Instruction fine-tuning
Queries 150 100
Epochs 600 50
Batch size 5 5$\times$200
Learning rate 1e-4 8e-4
LR schedule Cosine Cosine
Optimizer AdamW AdamW
Contrastive language-scene pre-training.
The pre-training is conducted on Grounded Scene Caption, ScanNet-200 detection data, ScanRefer, and Multi3DRef data unless otherwise specified. Following the prior detector settings [62], we initialize 150 queries via farthest point sampling (FPS) from the entire scene. The temperature for all contrastive learning activities is set at 0.1. CLASP is initialized using the pre-trained 3D detector Mask3D [62], incorporating similar mask losses as follows:
$\mathcal{L}_{mask}=\lambda_{SF}\mathcal{L}_{SF}+\lambda_{dice}\mathcal{L}_{dice},$ (4)
where $\mathcal{L}_{SF}$ represents the sigmoid focal loss, and $\mathcal{L}_{dice}$ denotes the Dice loss for segmentation. We set $\lambda_{BCE}=5$ and $\lambda_{dice}=2$. The overall CLASP loss is defined as:
$\mathcal{L}_{CLASP}=\mathcal{L}_{mask}(S_{mask},T_{mask})+\lambda_{cls}\mathcal{L}_{cls}(S_{text},T_{text}),$ (5)
where $S_{mask}$ represents the similarity matrix between instances and points, and $S_{text}$ is the similarity matrix between text phrases and instances. $T_{mask}$ and $T_{text}$ are the corresponding ground-truth mappings. Note that $T_{mask}$ is computed using Hungarian matching. $\mathcal{L}_{cls}$ represents the sigmoid focal loss. We omit the background class because it does not correspond to a well-defined text prompt. Further implementation details can be found in Mask3D [62].
Multi-task instruction fine-tuning.
During instruction fine-tuning, the LLM is frozen, except for the projection and language heads. We utilize LoRA [32] to fine-tune the LLM. Grounded 3D-LLM is initialized with a pre-trained CLASP network. To minimize token costs, only 100 queries are utilized. The instruction fine-tuning dataset includes data conversation using various task-specific instruction templates as detailed in Sec. G. The converted datasets comprises ScanRefer, Multi3DRef, ScanNet-200 detection, along with Scan2Cap and ScanQA dataset. During training, a visual query is considered a positive match only if its mask IoU with the ground-truth object is greater than $0.3$. An additional projection layer is employed to align these features with the language embedding space of the LLM. The overall language loss is used as follows:
$\mathcal{L}_{LLM}=\mathcal{L}_{\text{language}}+\mathcal{L}_{ref}(S_{ref},T_{ref}),$ (6)
where $L_{\text{language}}$ is the default cross-entropy loss as LLAMA [64] and $L_{ref}$ computes the sigmoid focal loss between the similarity matrix of referent tokens $S_{ref}$ and the ground-truth referent-instance mapping $T_{ref}$.
Appendix C Additional Ablation Studies
Designs of cross-modal interactor. As illustrated in Table A-8, we evaluate the architectures of the scene-text interaction network, which includes cross-modal interactions using bidirectional cross attentions [49] subsequent to the BERT encoder and the prevalent Q-Former structure [43]. The results show that incorporating bidirectional cross-attention modules in cross-modal interactor enhances performance on grounding metrics and dense captioning metrics. This enhancement indicates that the module derives greater benefits from CLASP pre-training, thereby establishing it as the default cross-modal interactor.
Table 8: Comparison of cross-modal interactors. “Bi-Direct C. A.” denotes the use of bidirectional cross-attention modules (default in the cross-modal interactor) for interacting with scene queries following a BERT encoder. “Q-Former” resembles BLIP-2 [43, 17] but incorporates our scene queries.
Cross- interactor types ScanRefer Scan2Cap
Acc@0.25 Acc@0.5 BLEU-4@0.5 CIDEr@0.5
Q-Former 47.2 44.0 34.9 70.4
Bi-Direct C. A. 47.9 44.1 35.5 70.6
Diversity of instruction templates. Typical computer vision tasks such as language grounding, object detection, and dense captioning require instruction templates to better adapt to pre-trained weights of large language models. We assess the impact of diverse instruction templates on language tasks in Table A-9. Results show that diverse instruction templates consistently yield significant improvements in tasks like dense captioning and 3D QA. Notably, in template-based captioning tasks (Scan2Cap), both CIDEr@0.25 and CIDEr@0.5 see improvements exceeding $8$ points. Employing a single instruction template also leads to unstable results in repeated experiments if diverse templates are not used. Therefore, the diversity of instruction-tuning templates is crucial for sustaining language capabilities in visual tasks.
Table 9: Comparison with and without using diverse instruction templates.
Diverse Templates Scan2Cap ScanQA
B-4@0.25 C@0.25 B-4@0.5 C@0.5 B-4 C
31.5 65.8 29.2 62.1 11.2 70.8
37.2 74.4 35.5 70.6 13.4 72.7
Effects of LLM model sizes. As shown in Table A-10, we compare the performance of GPT-2, Tiny-Vicuna-1B, Vicuna-7B, and Vicuna-13B. Smaller models like GPT-2 struggle with soft referent token learning. Conversely, Grounded 3D-LLM models, ranging from 1B to 13B parameters, greatly outperform GPT-2 and show marginal improvements in dense captioning tasks, yet they experience a slight decline in 3D visual QA performance. Notably, even when only the projection layer is trainable (without LoRA), Grounded 3D-LLM (Vicuna-7B) still achieves comparable performance. Overall, these models do not significantly excel in language tasks, suggesting that models up to Vicuna-1B are adequate for existing grounding benchmarks, indicating that the current task scope may not fully challenge larger-scale LLMs.
Table 10: Comparison of different LLMs. For GPT-2, we train the overall backbones with a smaller language head. In models without LoRA, only the projection layers are trained.
LLM LoRA Scan2Cap ScanQA Trainable Params
B-4@0.25 C@0.25 B-4@0.5 C@0.5 B-4 C
GPT-2 25.1 36.7 24.2 35.2 4.91 41.9 169 MB
Tiny-Vicuna-1B 37.0 74.4 35.5 70.6 13.4 72.7 157 MB
Vicuna-7B 38.2 75.6 36.0 71.2 12.3 71.8 364 MB
Vicuna-7B 38.1 74.3 35.1 71.0 11.8 70.6 330 MB
Vicuna-13B 39.5 75.4 36.1 69.2 12.1 71.6 487 MB
Appendix D Result Visualizations
Result visualization. We visualize the results in Fig. A-4 for a single model Grounded 3D-LLM that solves all downstream tasks, as opposed to per-task fine-tuning on downstream tasks[88, 31, 11].
Failure cases. As depicted in Fig. A-5, the primary reasons for our model’s failures in the grounding task include missing grounding objects (a, c), excessive grounding of semantically incorrect objects (b), and errors in detection/segmentation (c).
Figure 4: One model to solve all tasks. Result visualization in multiple downstream benchmarks for Grounded 3D-LLM. Please zoom in for better visualization.
Figure 5: Failure cases. Please zoom in for better visualization.
Appendix E Grounded Scene Caption Dataset
Data generation details. For the creation of Grounded Scene Captions (G-SceneCap), we selected the ScanNet-200 dataset due to its extensive range of semantic object categories. The data generation process employs heuristics to ensure that each local scene caption is concise, with fewer than 256 words. Initially, we identify each anchor object within a 2-meter radius to tally the objects. If this count exceeds 15, we reduce the radius to maintain this limit. Subsequent object selection is conducted through random sampling, with selection probabilities ranging from 0.6 to 0.9. This method prevents the selection of repetitive object types while preserving the anchor. Finally, for each caption, we generate program-derived relations that use existing objects as the anchor following SR3D [2]. These relations are then integrated into the caption. The overall process always keeps both the anchor and target object IDs.
Post processing. As ChatGPT/GPT-4 may generate errors such as incorrect output formats or erroneous object IDs, we filter out these wrong captions from the dataset.
Dataset language components visualization. For the grounded scene caption dataset, we visualize in Fig. A-6 the relationships between noun-adjective pairs and spatial preposition-noun pairs across all languages, along with a word cloud of all nouns, to demonstrate the diversity of our curated dataset.
Figure 6: Data statistics of grounded scene caption. Left: Diagram of spatial preposition-noun statistics, Middle: Noun-adjective statistics diagram, Right: Noun word cloud.
Dataset visualization. We provide additional visualizations of grounded scene caption examples in Fig. A-7.
Figure 7: Example visualization of grounded scene caption dataset. The “[ ]” symbol marks each grounded phrase. The color of the phrase corresponds to the segmented point cloud.
Appendix F Data Generation Prompts
Fig. A-8 displays the object caption prompts, including the VLM captioning prompt [70] and the subsequent condensation of the caption into a descriptive phrase using ChatGPT. These object captions are then compiled to create local scene captions, as shown in Fig. A-9, while retaining their object IDs. Finally, spatial relationships derived from program-generated code are incorporated using the prompt depicted in Fig. A-10.
Figure 8: Object caption prompt for VLM and ChatGPT (used in step 1 of data generation). The above is the prompt for CogVLM that utilizes the object labels from ScanNet-200. Below is the prompt for ChatGPT input, where “System” denotes the system prompt and “User” denotes user input.
Figure 9: GPT-4 prompt for scene caption (used in step 2 of data generation). “System” denotes the system prompt and “User” denotes user input for GPT-4.
Figure 10: GPT-4 prompt of including rule-based relation (used in step 3 of data generation). “System” denotes the system prompt and “User” denotes user input for GPT-4.
Appendix G Task-specific Instruction-tuning Templates
Grounded 3D-LLM simultaneously undertakes vision tasks—such as 3D visual grounding, 3D object detection—and language tasks, like question answering. Each task demands specific, diverse templates to enable the LLM to naturally provide scene-aware answers tailored to particular dataset specifications. We show several instruction templates in the following figures.
Object detection. The detection tasks employ various question templates for each category, along with three answer templates to accommodate various detection outcomes – single, multiple, or no objects detected – as illustrated in Fig. A-11.
Figure 11: Instruction templates for object detection.
Single and multi-object grounding. The language grounding’s question template, as depicted in Fig. A-12, directs the model to identify referred objects. For the ScanRefer grounding dataset, the answer templates indicate single-object grounding. The answer templates for the Multi3DRef dataset accommodate various grounding scenarios—single, multiple, or no grounded objects.
Figure 12: Instruction templates for single- or multi- object grounding.
Question answering. For ScanQA, we append suffixes to brief outputs (Fig. A-13) to indicate the outputs with only phrases or word output as the dataset annotation guideline.
Figure 13: Instruction templates for ScanQA.
Figure 14: Instruction templates for Scan2Cap.
Dense captioning. Question templates for Scan2Cap require the model to describe object appearance followed by spatial relations, as illustrated in Fig. A-14.
Appendix H Extension to embodied dialogue and embodied planning
Figure 15: GPT-4 Prompt of embodied dialogue. “System” denotes the system prompt and “User” denotes user input for GPT-4. “{Grounded scene caption}” is filled with our grounded scene caption data.
Figure 16: GPT-4 Prompt of embodied planning. “System” denotes the system prompt and “User” denotes user input for GPT-4. “{Grounded scene caption}” is filled with our grounded scene caption data.
To demonstrate potential robot applications in embodied dialogue and planning, we trained a model using a prior instruction-tuning dataset combined with a generated embodied dialogue and planning dataset. Fig A-15 and A-16 list the prompts for grounded scene caption-based dialogue and planning. The training dataset includes 11.9K dialogue examples and 4.4K planning examples. Training setups are the same as Sec.4.1.
Result visualization.
As there is no well-defined benchmark for evaluating both phrase grounding and language understanding, we visualize several results for embodied dialogue (Fig.17) and embodied planning (Fig.18). These examples show the potential of Grounded 3D-LLM’s grounding capability to enhance robot navigation and manipulation.
Figure 17: Result visualization for embodied dialogue. Please zoom in for better visualization.
Figure 18: Result visualization for embodied planning. Please zoom in for better visualization.