### v2

# Grounded 3D-LLM with Referent Tokens

Abstract
Prior studies on 3D scene understanding have primarily developed specialized models for specific tasks or required task-specific fine-tuning. In this study, we propose Grounded 3D-LLM, which explores the potential of 3D large multi-modal models (3D LMMs) to consolidate various 3D vision tasks within a unified generative framework. The model uses scene referent tokens as special noun phrases to reference 3D scenes, enabling it to handle sequences that interleave 3D and textual data. Per-task instruction-following templates are employed to ensure natural and diversity in translating 3D vision tasks into language formats. To facilitate the use of referent tokens in subsequent language modeling, we provide a large-scale, automatically curated grounded scene-text dataset with over 1 million phrase-to-region correspondences and introduce Contrastive Language-Scene Pre-training (CLASP) to perform phrase-level scene-text alignment using this data. Our comprehensive evaluation covers open-ended tasks like dense captioning and 3D question answering, alongside close-ended tasks such as object detection and language grounding. Experiments across multiple 3D benchmarks reveal the leading performance and the broad applicability of Grounded 3D-LLM. Code and datasets are available at the project page.
Index Terms:
Large Multi-modal Model, 3D Scene Understanding
1 Introduction
The pursuit of a unified model for 3D scene understanding, capable of performing various vision tasks, is a longstanding and significant challenge. Existing 3D scene-level models are typically customized for specific downstream tasks, including 3D object detection [1, 2, 3, 4], language grounding [5, 6, 7, 8, 9], question answering [10, 11, 12, 13], and dense captioning [14, 15, 16, 8, 17, 18, 13]. Recent 3D large multi-modal models (LMM), such as Chat-3D [19] and LL3DA [20], facilitate visual interactions based on detected objects for text reasoning, yet their capabilities remain confined to text-level tasks. 3D-LLM [21], which extends 2D VLMs [22] to 3D, grapples with challenges in 3D spatial reasoning due to the intrinsic knowledge derived from perspective views. This results in suboptimal performance in crucial localization tasks like object localization and language grounding, thereby limiting its practical applicability. The capacity of large multi-modal models (LMMs) to pinpoint objects or regions in complex environments is vital, especially for applications such as VR/AR, robotics, interactive embodied agents, object navigation, and manipulation.
However, the exploration of a unified model for 3D perception and reasoning remains scarce, leaving it as an open problem. In our study, by leveraging the generative modeling capability of large language models (LLM), we introduce Grounded 3D-LLM for integration of multiple 3D vision tasks in linguistic formats (Fig. 1).
Its model design is inspired by the natural text, ”Two brown chairs are near the white nightstand,” which contains two noun phrases: ”two brown chairs” (region) and ”the white nightstand” (singular object). We refer to their corresponding 3D objects in the scene as ”phraseable objects,” indicating that each can be described in a text phrase and that together they are interrelated within the contextual description. Without loss of generality, we aim to represent these phraseable objects using a special token, termed the ”referent token” <ref>. Each referent token supports decoding to identify the corresponding objects (singular or plural) in a 3D scene. In the above example, <ref> can refer to items like “the white nightstand” or “two brown chairs”, substituting the corresponding text tokens in the language. The use of referent tokens as manageable noun phrases in linguistic descriptions offers two advantages: (i) a referent token can serve as either an alternative to or a companion for text phrases without altering the contextual description, as depicted in Fig. 1. and (ii) it enables the localization of nouns or text phrases within the language. Thus, this design positions referents as a unified interface, facilitating 3D grounding capabilities across various 3D vision tasks.
To accomplish this, the language modeling of the referent involves two high-level steps: (i) A more granular, phrase-level contrastive pre-training of the vision-language encoder in a large-scale scene-text dataset rich in phrase-level correspondence, and (ii) referent-aware instruction-following fine-tuning, which associates referent tokens with their respective object embeddings.
Figure 1: Unifying scene-level 3D vision tasks via task-specific prompts in Grounded 3D-LLM framework. The interleaved referent token <ref> in language modeling allows the identification of external scene referents through masked decoding in raw point clouds. Words within <p> and </p> denote the accompanying text phrases. Matching colors are shown in the figure to connect phrases with their respective object masks.
Specifically, unlike CLIP, which aligns text with the whole image, we first perform contrastive alignment between text primitives – noun phrases in the contextual description – with their corresponding visual embeddings from 3D scene encoder. This scene-level pre-training process is referred to as Contrastive LAnguage-Scene Pre-training (CLASP). However, existing paired scene-text data at the phrase level (referred to as grounded scene-text data in this paper) remain scarce. To address this, we develop an automated pipeline to synthesize grounded scene-text data using a densely annotated 3D detection dataset. The resulting dataset, called Grounded Scene Caption (G-SceneCap), comprises 107K scene captions with over 1+ million phrase-level correspondence. Compared to existing scene-text datasets, it provides richer phrase-level correspondence, averaging one object every nine tokens, as shown in Table II. Besides, existing human-annotated 3D vision-language datasets [5, 23] are also into grounded language forms.
After the phrase-level scene-text alignment, we conduct multi-task instruction fine-tuning for Grounded 3D-LLM to associate the referents with their respective phraseable objects. We convert various existing grounded scene-text datasets into instruction-following format, where referent tokens are interleaved as alternatives or companions to the text phrases (examples shown in Fig. 1). During training, these special referent tokens are supervised to decode into specified object embeddings, identifying their external scene referents. These task-specific instruction-following conversions encompass a range of existing 3D vision-language tasks – from single to multi-object grounding, instance segmentation (object detection), to 3D QA and dense captioning. We compare this model’s prompt support and task versatility with previous large multi-modal models in Table I. We believe this represents a significant step toward unifying various 3D vision tasks within generative language modeling.
To summarize, our contributions are as follows:
  • We propose Grounded 3D-LLM, which establishes correspondence between 3D scenes and natural language using referent tokens. This method facilitates scene referencing and effectively models various 3D vision-language problems within a unified language modeling framework.
  • We first provide an automated, curated grounded scene caption dataset with over 1 million phrase-level correspondences. Experiments show that CLASP, using this data in both supervised and zero-shot text settings, demonstrates broad generalization in phrase-level grounding.
  • Without requiring specialized models or task-specific fine-tuning, the single model Grounded 3D-LLM achieves top-tier performance in most downstream tasks among generative models, particularly in grounding problems.
TABLE I: Comparison of 3D multi-modal models. We refer to instance segmentation as Inst.Seg., object box detection as Obj.Det., single-object grounding as Grd., point-level grounding as Point-Grd., multi-object grounding as Multi-Obj Grd., question answering as QA., and dense captioning as Cap..
Method LLM Prompts Tasks
Text Vision Inst.Seg. Obj.Det. Grd. Point-Grd. Multi-Obj Grd. QA. Cap.
PointGroup [1]
Mask3D [4]
Multi3DRef [23]
BUTD-DETR [24]
3D-VisTA [25]
Chat-3D [19]
Chat-3D v2 [26]
3D-LLM [21]
LL3DA [20]
Grounded 3D-LLM
2 Related Works
2D large multi-modal models. Vision-language models [27, 22] are proposed to bridge visual and textual models via image-level contrastive learning. Their direct application enables open-world understanding [28, 29]. GLIP and its subsequent works [30, 31, 32] have innovatively formulated object detection as grounding problems, leveraging detection or pseudo-labels to align semantics. With stronger language models, various powerful large multi-modal models are capable of handling tasks including image captioning [33, 34], VQA [35, 36, 37, 38], and object detection/segmentation [32, 39]. Besides, incorporating multi-modal tokens into language modeling has been validated, including visual understanding [34, 40, 41, 42], robot control [43], and interleaved text-image content generation [44, 45]. On the other hand, Chen et al. [46] demonstrates that existing 2D VLMs lack 3D spatial reasoning, primarily due to insufficient 3D spatial data.
3D scene understanding. In 3D scene understanding, prior works [47, 48, 49] have advanced large-scale unsupervised learning for foundational models. Language is increasingly used to encapsulate user intentions due to its broad applicability, ranging from closed-set object detection [1, 2, 4, 50, 24] to open-set 3D detection [51, 52, 53, 54, 55, 56]. Typically, these 3D detectors lay the groundwork for subsequent scene-language tasks. Such tasks encompass 3D Visual Grounding [5, 57, 7, 23, 9, 58, 6, 59], where models pinpoint objects within a 3D scene based on language queries. They also include 3D Question Answering [10, 60, 61], which demands robust spatial perception and reasoning. Another task, 3D Dense Captioning [14, 62, 63], involves localizing and describing all scene objects, necessitating a comprehensive understanding of object locations and attributes. While initial efforts were often task-specific and lacked generalizability, some methodologies [17, 8] have integrated 3D visual grounding and dense captioning tasks, leveraging their complementary nature. Recent endeavors such as 3D-VisTA [25] and 3D-VLP [13] aim to establish a universal 3D visual-language framework by pre-aligning 3D scenes with language. In contrast, most approaches develop specialized models or require per-task training. They also rely on off-the-shelf 3D detectors.
3D large multi-modal models. 3D object-level LMMs [64, 65, 66, 67] have effectively bridged the gap between 3D visuals and texts by leveraging extensive 3D object datasets [68]. This has advanced the development of 3D object-level LMM [69, 70, 21, 71, 72, 67, 73]. However, these models struggle with interpreting complex spatial interrelationships in 3D scenes. For scene-level LLMs, models such as Chat-3D [19], Scene-LLM [74], LL3DA [20], and LEO [75] enable scene-level dialogue, achieving notable effectiveness in tasks such as question answering and captioning. Chat-3D v2 [26] enables language grounding by using object identifiers for effective referencing. 3D-LLM [21] improves scene understanding by incorporating positional embeddings and location tokens. Our approach utilizes special referent tokens for flexible scene modeling and supports diverse tasks including object detection and grounding. Additionally, our model functions as a generalist model without requiring task-specific fine-tuning.
3 Method
This section consists of three parts. Sec. 3.1 examines vision-language alignment at three levels. Sec. 3.2 details the Grounded 3D-LLM, from vision-language pre-training to multi-task instruction fine-tuning with referent tokens. Finally, Sec.3.3 describes the construction of the large-scale grounded scene-text dataset for pre-training, while Sec.3.4 illustrates the conversion of existing scene-text data to instruction-following formats for LLM fine-tuning.
3.1 Preliminary: Vision-Language Correspondence
The significance of vision-language alignment in 2D foundation models, exemplified by the Contrastive Language-Image Pre-training (CLIP) [27], is well-established. We explore three levels of vision-language alignment:
  • Sentence-level Correspondence: CLIP models [27, 22] introduced contrastive learning between image and sentence-level text embeddings. However, the limited size ($\sim$100K-scale) of paired scene-text datasets[5, 57, 23] hinders broader application, and the alignment lacks object-level grounding.
  • Word-level Correspondence: Initially introduced by GLIPs [30], it reinterprets object detection labels as phrase grounding tasks. However, this reliance on existing detection labels does not adequately capture appearance and spatial relationships.
  • Phrase-level Correspondence: As referred to as phrase-to-region correspondence in prior work [76], this constraint requires linking text phrases to specific objects (singular or plural) within a contextual description including category, attributes and etc.. However, such 3D datasets remain scarce, especially in 3D scenes.
Figure 2: Main pipeline for Grounded 3D-LLM. The training process for Grounded 3D-LLM is divided into two key steps. Firstly, CLASP utilizes an extensive grounded scene-text dataset (at the phrase level) to pre-train a 3D point cloud encoder and a cross-modal interactor. The subsequent step involves multi-task instruction fine-tuning, which interlaces referent tokens within the instructions and responses, thereby facilitating flexible 3D scene understanding tasks.
Our motivation. To enable the natural referencing of scene objects in language modeling, we aim to model phrase-level correspondences, allowing phraseable objects to be represented as referent tokens. In other words, referent tokens can substitute for or accompany noun phrases without altering the language structure. Inspired by Llava [35], we first adopt contrastive pre-training for the vision-language encoder at the phrase level using grounded scene-text data. During instruction-following fine-tuning, the LLM associates the newly proposed referent token with these phraseable objects. For example, in the phrase ”two brown chairs near the table,” the average word embeddings of ”two brown chairs” align with the 3D visual embeddings of the corresponding chairs. By inputting these text-aligned object embeddings, the LLM can reference these phraseable objects using the special referent. This can be formatted as <p>two brown chairs</p> <ref> near the table or simply as <p>object</p> <ref>. The token pairs <p> and </p> indicate phrase boundaries [39, 42].
3.2 Grounded 3D-LLM
Based on the aforementioned motivation, the main pipeline of Grounded 3D-LLM is illustrated in Fig. 2. To reference phraseable objects in language modeling, a vision-language model is first pre-trained on large-scale grounded scene-text data to align text phrases with their corresponding 3D objects. Subsequently, a large language model (LLM) is fine-tuned using multi-modal instruction-following data, where referent tokens serve as interleaved “soft” text tokens representing the phraseable objects. Per-task instruction-following templates are employed to address the diverse range of 3D vision tasks within the unified language modeling framework.
Step 1: Contrastive language-scene pre-training (CLASP). In the same spirit as CLIP [27], CLASP is pre-trained to align 3D scene embeddings with textual descriptions using grounded scene-text data. Each paired grounded scene-text annotation contains phrase-to-region correspondences, linking noun phrases inside the text and their respective 3D object IDs (singular or plural). CLASP co-trains a 3D encoder, a text encoder, a set of object queries, and a cross-modal interactor (CMI), as shown on the left side of Fig. 2. Specifically, a sparse convolutional network [77] serves as the 3D multi-level feature encoder, extracting point-wise embeddings $P$ as a $K$-level feature pyramid, while BERT [78] extracts text queries $T$. For each pyramid level, a cross-modal interactor (CMI) is used to extract object embeddings conditioned on both point cloud embeddings and their respective text descriptions. Specifically, CMI interacts with $N$ learnable object queries $Q$ as proxies to connect an $M$-word text (extracted from BERT) with point clouds. Initially, these object queries interact with $k$-level point-wise embeddings via cross-attention. Then, as shown in Fig. 3, to incorporate the contextual text description, they interact with text embeddings through plain one-way cross-attention or bi-directional cross-attention [32].
Figure 3: Designs of Cross-Modal Interactor (CMI). Sub-figure (b) shows the bi-directional cross-attention design inspired by GroundingDINO [32], which is the default choice.
Figure 4: Grounded Scene Caption dataset generation pipeline. This pipeline leverages privileged knowledge in 3D, including dense object labels, multi-view images, and global spatial relationships, along with the capabilities of ChatGPT tools and spatial relationship programming tools, to establish accurate scene-level captions with grounding information.
As shown in Fig. 2 (left), CMI yields both object and text embeddings, followed by a small head that transforms them into the textual embedding space: $P_{text}$ for positive text embeddings and $Q_{text}$ for object embeddings, as well as the mask embedding space: $P_{mask}$ and $Q_{mask}$ for mask embeddings of both, respectively:
$\left[P_{mask},Q_{mask}\right],\left[Q_{text},P_{text}\right]=\text{CMI}(P,T,Q).$ (1)
As depicted in Fig. 2, we compute the alignment scores $S_{mask}$ for per-point mask classification and $S_{text}$ for the textual embedding similarities between noun phrases and object queries:
$S_{mask}=P_{mask}Q_{mask}^{\top},\quad S_{text}=P_{text}Q_{text}^{\top},$ (2)
where we ignore the temperature $\eta=0.1$. The per-point mask classification for $S_{mask}$ is computed to classify each point against the object queries using bipartite matching [79, 80]. Another contrastive loss for noun phrases and object queries, $S_{text}$, is supervised by the phrase-to-region correspondence. During inference, the mapping from text phrases to point-level masks can be computed using both similarity matrices, e.g., by employing the top-scored indices as the mapping.
Step 2: Instruction fine-tuning with referent tokens. During instruction-following fine-tuning, Grounded 3D-LLM models input and output in a unified multimodal language format, fine-tuning a LLM based on the pre-trained CLASP. Specifically, the LLM inputs scene encoding (all object query embeddings) from the frozen CLASP with a plain detection prompt [30]. It aims to learn to use the referent token <ref> to reference the scene, given a large amount of grounded instruction-following data.
Given a single grounded instruction-following data point, the text is interleaved with referent and each referent has its mapping to some phrasesable objects (referent correspondence). We first convert the instruction and response into the USER-ASSISTANT format before feeding them into the LLM. An example is shown below:
[All object embeddings] USER: Describe the
object <ref> in the scene.
ASSISTANT: <p> A black chair </p> <ref>
with four legs.
To enable closed-loop dialogue with the LLM using scene referents, we model referents as their corresponding objects in both language input and output. Specifically, for language input, the referents (visible to the model) are directly replaced with their respective object query embeddings, followed by a linear projection to match the LLM channel size. For language output, the referents serve as placeholders, with their hidden embeddings $R_{text}$ aligned with the corresponding object query embeddings. During fine-tuning, we compute the similarity matrix between the referent and all object query embeddings $Q_{text}$:
$S_{ref}=R_{text}Q_{text}^{\top}.$ (3)
The contrastive loss for $S_{ref}$ is supervised with the referent correspondence and applied with the same temperature $\eta$ as CLASP. During inference, the output referents are decoded with a decoder (two MLP layers) to extract hidden embeddings $R_{text}$, retrieve the respective object queries, and obtain their associated masks through prior CLASP.
Besides, we ensure linguistic naturalness by formatting referents with accompanying text phrases, such as “<p> three nearby chairs </p> <ref>.”. If no category is provided in scene-text annotations, we use the generic term “object” instead. For tasks requiring grounding or localization, we append “(with grounding)” to instructions to enable referencing in the language output.
Discussion: Referent token as localization token. Prior work, 3D-LLM [21], uses discrete location tokens, limiting localization accuracy and hindering grounded visual chat tasks. Chat-3D v2 [26] pairs object identifiers with features as text inputs but incurs high token costs and relies on off-the-shelf detectors and per-object feature extractors [67]. In contrast, our model jointly extracts per-object features through large-scale, fine-grained scene-level pre-training of CLASP. The design of the Referent token allows decoding into multiple instances, reducing token costs and increasing localization efficiency. Additionally, it is the first to support point-level mask decoding.
3.3 Generating Grounded Scene-Text Data for Pre-training
In this subsection, we describe two approaches to generate grounded scene-text data for pre-training: (i) automatic generation of grounded scene captions, and (ii) transformation of existing language data into grounded formats.
Figure 5: Grounded Scene Caption visualization. In each grounded text, a phrase-to-region correspondence is formatted as “[object_phrase object_IDs]”, where “object_phrase” refers to a noun phrase and “object_IDs” denotes the corresponding object IDs in the groud-truth annotation. The bottom line demonstrates explicit phrase-to-region correspondence example between the noun phrases (character positions) and the corresponding objects.
TABLE II: Grounded scene-text dataset comparison. “Sent.”, “Obj.”, “Multi.Obj.”, and “Corr.” refer to “Sentence”, “Object”, “Multi-objects”, and “Correspondence”. Tokens are calculated using BERT [78]. For sentences with single-object correspondence, each text is paired with one instance.
Dataset #3D Scan #Tokens #Text #Tokens/Text Correspondence
#Corr. Type Corr/Token
ScanRefer [5] 0.7K 1.17M 52K 22.6 52K Sent.-Obj. 4.4%
Sr3D [57] 1.5K 1.10M 84K 13.1 84K Sent.-Obj. 7.7%
Sr3D+ [57] 1.5K 1.48M 115K 12.9 115K Sent.-Obj. 7.7%
Nr3D [57] 0.7K 0.617M 42K 14.7 42K Sent.-Obj. 6.8%
Multi3DRefer [23] 0.7K 1.20M 62K 19.4 87K Sent.-Mult.Obj. 7.2%
ScanScribe [25] 3.0K 18.5M 278K 66.5 278K Sent.-Obj. 1.5%
Grounded Scene Caption 1.5K 9.19M 107K 85.9 1.02M Phrase-Multi.Obj. 11.1%
Grounded Scene Caption dataset generation. The dataset generation for the Grounded Scene Caption dataset (G-SceneCap) leverages ChatGPT and 2D vision-language models, utilizing dense object annotations from existing 3D scan datasets [81], as illustrated in Fig. 4:
Step 1: Bootstrapping object captions with GT label correction. Using 3D real-scan datasets, we annotate each object with the vision-language model CogVLM [38], using the images of the largest visible areas. Inconsistent annotations are rectified using raw instance labels.
Step 2: Condensing objects in local scenes into a caption. For each enumerated anchor object, we form an initial object set by randomly selecting a group of nearby objects. Their captions and coordinates $(x,y,z)$ are input into GPT-4 for captioning, which requires referencing object phrases with their object IDs in the format “[object_phrase object_ID]” in the caption.
Step 3: Adding rule-based relations. To enrich scene captions, we integrate program-generated spatial relationships from Sr3D [57]. By selecting an anchor object from the set in step 2, we apply the spatial relation rules (e.g., between, supporting, nearest, back) to include related objects. GPT-4 then combines these relationships into the prior caption from step 2.
A specific coarse-to-fine scene caption generation example with phrase-to-region correspondence, including step-by-step intermediate outputs, is illustrated in Fig. 10. The data generation prompts are detailed in the supplementary files.
Grounded conversion of existing scene-text data. In this work, we employ ChatGPT [82] to extract positive phrases for ScanRefer and Multi3DRef datasets for training and evaluation. We refer to the grounded datasets as G-ScanRefer and G-Multi3DRefer. The transformation of all detection labels into a fixed text prompt is called G-Detection as GLIP [30].
Figure 6: Converting grounded scene-text data to instruction-following format. This figure shows an instruction-following conversion example for the multi-object grounding task. Note that the referent correspondence is converted from the phrase-to-region correspondence of grounded scene-text annotation.
Comparsion with scene-text datasets. Llava [35] uses GPT-4 to generate multi-modal data based on 2D boxes. Similarly, 3D-LLM [21] relies only on 3D boxes to create instruction-following data. Other 3D works [53, 54] add cross-view information to capture view-level differences and interactions, improving open-vocabulary detection. ScanScribe [25] creates scene annotations from human scene graphs. Sceneverse [83] proposes generating scene graphs automatically, then rephrasing them into captions. Multi3DRef [23] generates multi-object correspondence using ScanRefer annotations.
Beyond these efforts, we focus on capturing noun relationships within natural language rather than generating object-level or region-level captions. Additionally, global-view spatial relationships are considered through GPT (Step 2) and rule-based programming (Step 3). As a result, our generated grounded scene-text data achieves the highest object density, averaging one correspondence per nine tokens, leading to over 1.02 million correspondences, as shown in Table II. This provides rich phrase-level correspondence in natural text and allows for seamless conversion to referent correspondence in the following subsection.
3.4 Converting to Grounded Instruction-Following Format for LLM fine-tuning
In this subsection, we describe the transformation of grounded scene-text data from Sec. 3.3 to the instruction-following format.
Converting 3D vision-language tasks to instruction-following formats. For existing 3D vision-language tasks, we can transform them into instruction-following formats. An example of conversion for multi-object grounding tasks including referent correspondence generation is shown in Fig. 6. These convertible 3D vision-language tasks include single and multi-object grounding, object detection, dense captioning, 3D QA, etc.. For each task, we utilize approximately 10-20 structured task-specific instruction-following templates. Please refer to the supplementary file for details on the task-specific instruction-following templates.
Extension to embodied planning and dialogue. To illustrate the potential applications of referent tokens in daily life and enable multi-round conversations, we prompt GPT-4 to generate embodied dialogue and planning based on the Grounded Scene Caption dataset, rather than using static templates. Unlike 3D-LLM [21], our approach produces not only plain dialogue text but also grounded instruction-following text with interleaved phrase-to-region correspondence (referent tokens). These dialogues adhere to several rules: they occur between a human and an embodied agent and involve tasks such as completing daily activities, discussing layouts, and conducting object-oriented QA based on the given caption. The embodied planning text must generate daily instructions and a step-by-step plan for interacting with objects within the scene caption. Refer to the supplementary files for the prompts used in specific dialogue data generation and the visualized results from our jointly trained model.
TABLE III: Evaluation of 3D scene-level LMMs as a generalist. Entries in gray denote models specialized for specific datasets. “Zero-shot” denotes models used directly without fine-tuning. “Specialist” and “Generalist” categorize models fine-tuned for particular tasks versus those trained jointly. Entries in grey denote the incorporation of ground-truth objects into the model inputs.
Method Generative Model ScanRefer Grd. Multi3DRef Grd. ScanQA Scan2Cap ScanNet-200 Det.
Acc@0.25
Acc@0.5
F1@0.25
F1@0.5
B-4 C
B-4@0.5
C@0.5
AP AP0.25 AP0.5
ScanRefer [5] 37.3 24.3
MVT [7] 40.8 33.3
3DVG-Trans [6] 45.9 34.5
ViL3DRel [9] 47.9 37.7
M3DRef-CLIP [23] 51.9 44.7 42.8 38.4
Scan2Cap [14] 22.4 35.2
ScanQA [10] 10.1 64.9
3D-VisTA [25] 50.6 45.8 13.1 72.9 34.0 66.9
Mask3D [5] 27.4 37.0 42.3
Zero-shot LMMs
LLM-Grounder [84] 17.1 5.3
Specialist LMMs
3D-LLM(Flamingo)
[21]
21.2 7.2 59.2
3D-LLM(BLIP2-flant5)
[21]
30.3 12.0 69.4
Chat-3D [19] 6.4 53.2
Chat-3D v2 [26] 35.9 30.4 7.3 77.1
LL3DA [20] 13.5 76.8 36.8 65.2
Generalist LMMs
LEO [75] 13.2 101.4 38.2 72.4
LL3DA [20] 13.3 75.7 36.0 62.9
Grounded 3D-LLM 48.6 44.0 44.7 40.8 13.2 75.9 35.0 70.2 12.1 16.8 18.9
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our model on the ScanRefer [5] (single-object grounding), Multi3DRef [23] (multi-object grounding), ScanNet-200 Detection [85] (object detection/instance segmentation), Scan2Cap [14] (dense object captioning), and ScanQA [10] (3D QA) benchmarks. These benchmarks are built on the ScanNet dataset [81], a large-scale indoor 3D scene dataset with 1,513 multi-modal scenes containing 3D point clouds, image sequences, and point-level instance segmentation annotations. The dataset is split into 1,201 scenes for training and 312 for validation, and all benchmarks follow these splits.
Metrics. We follow prior works and adopt the commonly used metrics for these benchmarks. For ScanRefer and Multi3DRef, we compute the Intersection over Union (IoU) at thresholds of $0.25$ and $0.5$. For ScanNet-200, we report 3D instance segmentation using average precision at $25\%$, $50\%$ overlap, and mean average precision (mAP) with IoU thresholds in the $[0.5\text{:}0.95\text{:}0.05]$ range across 200 classes. For Scan2Cap and ScanQA, we evaluate language generation with BLEU-4 and CIDEr, weighted by IoU above $0.25$ or $0.5$.
To maintain consistency with prior evaluations across the ScanRefer, Multi3DRef, and Scan2Cap benchmarks, we extract 3D bounding boxes by taking the minimum and maximum values from the predicted masks, despite our model producing point-level predictions. For Multi3DRef, we apply a score threshold of $0.3$ to filter the predicted objects.
TABLE IV: Effects of diverse instruction-following templates on language tasks.
Diverse Templates ScanQA Scan2Cap
B-4 C B-4@0.25 C@0.25 B-4@0.5 C@0.5
11.2 70.8 31.5 65.8 29.2 62.1
13.2 75.9 36.9 74.1 35.0 70.2
Implementation Details. Unless otherwise specified, CLASP pre-training is conducted on Grounded Scene Caption, ScanNet-200 detection data, ScanRefer, and Multi3DRef data. 150 object queries via farthest point sampling (FPS) is initialized for CLASP and initialized with pre-trained 3D detector Mask3D [4] to accelerate the training. We incorporate the same per-point mask classification loss $\mathcal{L}_{mask}$ with both sigmoid focal loss and Dice loss for segmentation. The overall loss function for CLASP is defined as:
$\mathcal{L}_{CLASP}=\mathcal{L}_{mask}(S_{mask},T_{mask})+\lambda_{cls}\mathcal{L}_{cls}(S_{text},T_{text}),$ (4)
where $T_{mask}$ is computed using Hungarian matching [4] and $T_{text}$ are the ground-truth mappings computed according to ground-truth phrase-to-region correspondence. $\mathcal{L}_{cls}$ represents the sigmoid focal loss.
During instruction fine-tuning, Grounded 3D-LLM adopts Tiny-Vicuna-1B as the LLM by default, keeping it frozen except for the intermediate projection and LoRA [86] layers. Its vision-language model is initialized with a pre-trained CLASP model. To minimize token costs, only 100 object queries are used. The fine-tuning leverages grounded instruction-following datasets, including ScanRefer, Multi3DRef, ScanNet-200 detection, Scan2Cap, ScanQA, scene captioning, embodied dialogue and embodied planning. An object query is considered a positive match only if its mask IoU with the ground-truth object exceeds $0.3$ otherwise the object will be ignored. The overall language loss is used as follows:
$\mathcal{L}_{LLM}=\mathcal{L}_{\text{lang}}+\mathcal{L}_{ref}(S_{ref},T_{ref}),$ (5)
where $L_{\text{lang}}$ is cross-entropy loss as LLAMA [87] and $L_{ref}$ computes the sigmoid focal loss between $S_{ref}$ and the ground-truth referent correspondence $T_{ref}$ from the grounded annotation.
TABLE V: Grounding ability of Grounded 3D-LLM. The term “all data” encompasses the G-ScanRefer, G-Multi3DRef, and G-SceneCap datasets used for pre-training.
# CLASP Pre-training data <ref> type ScanRefer Grd. Multi3DRef Grd.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5
(a) One-to-Many 6.5 5.8 6.4 5.3
(b) G-ScanRefer One-to-Many 43.2 36.6 40.2 36.9
(c) G-SceneCap One-to-Many 47.6 43.4 44.3 40.1
(d) All data One-to-One 46.5 43.3 44.2 40.7
(e) All data One-to-Many 47.9 44.1 45.2 40.6
TABLE VI: Comparison of CLASP in phrase grounding tasks. “Point-supervision” involves using point-level instance segmentation supervision. “End-to-end models” are those that can be trained either fully end-to-end or using separate detectors.
Method Point- supervision End-to -end ScanRefer Grd. Multi3DRef Grd. ScanNet200 Det.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5 AP AP@0.25 AP@0.5
BUTD-DETR [24] 52.2 39.8
ViL3DRel [9] 47.9 37.7
3D-VisTA [25] 50.6 45.8
M3DRef-CLIP [23] 51.9 44.7 42.8 38.4
Mask3D [4] 27.4 37.0 42.3
CLASP 53.2 46.7 51.5 47.3 27.2 36.4 42.1
TABLE VII: Effects of pre-training datasets for CLASP. $\dagger$ represents using ground-truth boxes.
# Pre-training Method Correspondence Level Training Data ScanRefer Grd. Multi3DRef Grd.
Acc@0.25 Acc@0.5 F1@0.25 F1@0.5
(a) M3DRef-CLIP Sentence Multi3DRef 55.0
(b) 3D-VisTA Sentence ScanScribe 41.1 36.4
(c) (b) w. ScanRefer 50.6 45.8
(d) CLASP Phrase G-SceneCap 46.2 35.9 39.3 36.1
(e) (d) w. G-ScanRefer 49.7 44.4 39.2 35.9
(f) (e) w. G-Detection 51.1 44.8 40.2 37.0
(g) (f) w. G-Multi3DRef 53.2 46.7 51.5 47.3
(h) (g) w/o. G-SceneCap 51.1 45.4 48.5 43.1
4.2 Main Results
Baselines. Existing baseline methods for downstream tasks are categorized into discriminative and generative approaches in Table III.
  • Discriminative models focus on single-task outputs with task-specific heads, covering single-object [9, 7, 6] and multi-object grounding [23]. 3D-VisTA [25] pre-trains a unified 3D backbone followed by task-specific fine-tuning, while Mask3D [4], a SOTA transformer-based 3D detector, is trained with instance segmentation heads.
  • Generative models include recent 3D LMMs [19, 26, 20, 21] without task heads, where specialist models require task-specific fine-tuning, and generalist 3D LMMs [20, 75] are trained in unified language format.
Result analysis. As shown in Table III, Grounded 3D-LLM, without task-specific fine-tuning, outperforms previous LLM-based models in most metrics, positioning it as a promising candidate for a unified framework in 3D scene understanding. Compared to prior 3D LMMs, it achieves the best grounding results, improving by over 10 points in both single and multiple object grounding. Additionally, it attains comparable outcomes in language-based tasks, even against LL3DA, which is limited to text output. Notably, it achieves leading performance with a CIDEr score of 75.9 and a BLEU-4 score of 13.2 in ScanQA. Existing generalist 3D LMMs, such as LEO [75] and LL3DA [20], support language tasks but are limited in localization. Furthermore, leveraging the one-to-many correspondence ability of the referent token, we report that Grounded 3D-LLM performs 3D instance segmentation, achieving 12.1 mAP. Despite its inferior performance compared to prior expert models like Mask3D, it demonstrates the potential to unify pure 3D vision tasks, with further improvements remaining as future work.
4.3 Ablation Studies
Diversity of task-specific instruction-following templates. Typical computer vision tasks such as language grounding, object detection, and dense captioning require instruction-following templates to better adapt to pre-trained weights of large language models. We assess the impact of diverse instruction-following templates on language tasks in Table IV. Results show that diverse instruction-following templates consistently yield significant improvements in tasks like dense captioning and 3D QA. Notably, in template-based captioning tasks (Scan2Cap), both CIDEr@0.25 and CIDEr@0.5 see improvements exceeding $8$ points. Employing a single instruction-following template also leads to unstable results in repeated experiments if diverse templates are not used. Therefore, the diversity of instruction-following templates is crucial for sustaining language capabilities in visual tasks.
Ablation studies of the referent token for grounding. As shown in Table V, comparing line (a) with others indicates that directly adapting existing closed-set detectors is ineffective for integrating referent tokens in Grounded 3D-LLM. Incorporating a grounded scene-text dataset for pre-training boosts performance to $43.2$ Acc@0.25 and $40.2$ F1@0.25. With only G-SceneCap for CLASP pre-training, Grounded 3D-LLM already achieves comparable results to all pre-training data.
Additionally, we analyze two types of referent tokens: one-to-one (d), where each object is matched with a unique token via Hungarian matching, and one-to-many (e), where a single token represents multiple objects. The latter generally yields better results.
Direct comparison of CLASP in phrase grounding tasks. CLASP performs phrase-level scene-text alignment for pre-training, allowing for direct comparison with previous grounding and detection methods. As shown in Table VI, CLASP excels in 3D grounding and detection benchmarks, particularly in ScanRefer grounding, Multi3DRef multi-object grounding, and ScanNet-200 detection, showcasing its enhanced phrase grounding capabilities compared to earlier models. Notably, in the Multi3DRef grounding task, CLASP outperforms specialized models by over 8 points, demonstrating its strong multi-object handling.
Dataset ablations for CLASP. Table VII further ablates the effects of datasets for the scene-text alignment:
(i) Zero-shot text evaluation. Without training on ScanRefer or Multi3DRef language data, our CLASP model (d) proves effective compared to previous models (a) and (b). Notably, despite 3D-VisTA’s data generation adhering to sentence-level grounding similar to ScanRefer, our model exceeds it by 4.9 points in Acc@0.25. Furthermore, in experiment (d), our direct evaluation on the Multi3DRef dataset demonstrates robust multi-object grounding ability.
(ii) Multi-dataset evaluation. We further investigate the effects of joint training with grounded scene-text datasets, progressively integrating G-ScanRefer (e), G-Detection (f), and G-Multi3DRef (g). This integration boosts performance, achieving 53.2 in Acc@0.25 on ScanRefer and 51.5 in F1@0.25 on Multi3DRef. Notably, Comparison between (g) and (h) shows that the G-SceneCap dataset consistently raises performance by +2.1 points in ScanRefer Acc@0.25 and +3.0 points in Multi3DRef F1@0.25, highlighting its effectiveness.
TABLE VIII: Comparison of cross-modal interactors (CMI).
Cross- interactor types ScanRefer Scan2Cap
Acc@0.25 Acc@0.5 B-4@0.5 C@0.5
One-way 47.2 44.0 34.9 70.4
Bi-directional. 47.9 44.1 35.5 70.6
Comparison of cross-modal interactors. Table VIII ablates two key multi-modal interaction components—cross-modal interactors of Fig. 3: text-to-object one-way CMI and bidirectional CMI. The results indicate that incorporating bidirectional cross-attention modules in cross-modal interactors enhances grounding performance (with a 0.7 improvement in Acc@0.25 on ScanRefer and a 0.6 improvement in BLEU-4@0.5 on Scan2Cap) and dense captioning metrics. This improvement suggests that deeper interactions between modalities provide greater benefits for CLASP.
5 Conclusion
We explore the potential of large multi-modal 3D models to unify various downstream 3D vision tasks in unified language modeling. By interpreting scene referents as special language tokens, Grounded 3D-LLM connects scene objects or regions with language, offering a natural way to localize noun phrases in the 3D environment. We demonstrate its general effectiveness across various 3D vision tasks; however, it shows limitations in performance compared to previous expert models, highlighting the disparity between specialist and generalist approaches. Scaling scene-text data to bridge this gap is an objective for future work.
Supplementary Files for Grounded 3D-LLM
The supplementary materials are organized as follows:
  1. 1.
    Additional ablation studies are included in Sec. 6.
  2. 2.
    Hyper-parameters of training are included in Sec. 7.
  3. 3.
    Visualizations of results and failure cases for downstream tasks are presented in Section 8.
  4. 4.
    Details on grounded scene-text data generation, visualizations of dataset statistics, and data generation prompts, are illustrated in Sec. 9.
  5. 5.
    Multi-task instruction-following templates are described in Sec. 10.
  6. 6.
    Extension to embodied dialogue and planning data is shown in Sec. 11
6 Additional Ablation Studies
Effects of LLM model sizes. As shown in Table A-IX, we compare the performance of GPT-2, Tiny-Vicuna-1B, Vicuna-7B, and Vicuna-13B. Smaller models like GPT-2 struggle with soft referent token learning. In contrast, Grounded 3D-LLM models, ranging from 1B to 13B parameters, significantly outperform GPT-2 and show marginal improvements on Scan2Cap, though they experience a slight decline on ScanQA. Notably, even with only the projection layer trainable (without LoRA), Grounded 3D-LLM (Vicuna-7B) achieves comparable performance, suggesting that the input scene embeddings are semantically rich and can be extracted by the LLM without intermediate trainable layers. Furthermore, the lack of improvement in language tasks with larger models indicates the current task scope may not fully challenge larger-scale LLMs.
TABLE IX: Comparison of different LLMs. For GPT-2, we train the overall backbones with a smaller language head. In models without LoRA, only the projection layers are trained.
LLM LoRA ScanQA Scan2Cap Trainable Params
B-4 C B-4@0.25 C@0.25 B-4@0.5 C@0.5
GPT-2 8.42 44.6 28.4 45.2 26.7 41.9 169 MB
Tiny-Vicuna-1B 13.2 75.9 36.9 74.1 35.0 70.2 157 MB
Vicuna-7B 13.3 74.1 37.8 74.3 35.4 70.8 364 MB
Vicuna-7B 12.4 70.9 37.4 74.2 34.7 69.9 330 MB
Vicuna-13B 12.9 73.2 38.5 74.6 35.9 70.4 487 MB
7 Implementation Details
The training hyper-parameters are detailed in Table A-X. All models are trained on 8 NVIDIA A100 (80G) GPUs.
TABLE X: Training hyper-parameters. The instruction fine-tuning batch size comprises 5 scenes, each containing up to 200 instructions.
Configuration CLASP pre-training Instruction fine-tuning
Queries 150 100
Epochs 600 50
Batch size 5 5$\times$200
Learning rate 1e-4 8e-4
LR schedule Cosine Cosine
Optimizer AdamW AdamW
8 Result Visualizations
Result visualization. We visualize the results in Fig. A-7 for a single model Grounded 3D-LLM that solves all downstream tasks, as opposed to per-task fine-tuning on downstream tasks[25, 21, 20].
Failure cases. As depicted in Fig. A-8, the primary reasons for our model’s failures in the grounding task include missing grounding objects (a, c), excessive grounding of semantically incorrect objects (b), and errors in detection/segmentation (c).
Figure 7: One model to solve all tasks. Result visualization in multiple downstream benchmarks for Grounded 3D-LLM. Please zoom in for better visualization.
Figure 8: Failure cases. Please zoom in for better visualization.
9 Grounded Scene Caption Dataset
Data generation details. For the creation of Grounded Scene Captions (G-SceneCap), we selected the ScanNet-200 dataset due to its extensive range of semantic object categories. The data generation process employs heuristics to ensure that each local scene caption is concise, with fewer than 256 words. Initially, we identify each anchor object within a 2-meter radius to tally the objects. If this count exceeds 15, we reduce the radius to maintain this limit. Subsequent object selection is conducted through random sampling, with selection probabilities ranging from 0.6 to 0.9. This method prevents the selection of repetitive object types while preserving the anchor. Finally, for each caption, we generate program-derived relations that use existing objects as the anchor following SR3D [57]. These relations are then integrated into the caption. The overall process always keeps both the anchor and target object IDs.
Post processing. As ChatGPT/GPT-4 may generate errors such as incorrect output formats or erroneous object IDs, we filter out these wrong captions from the dataset.
Dataset language components visualization. For the grounded scene caption dataset, we visualize in Fig. A-9 the relationships between noun-adjective pairs and spatial preposition-noun pairs across all languages, along with a word cloud of all nouns, to demonstrate the diversity of our curated dataset.
Figure 9: Data statistics of grounded scene caption. Left: Diagram of spatial preposition-noun statistics, Middle: Noun-adjective statistics diagram, Right: Noun word cloud.
Dataset visualization. We provide additional visualizations of grounded scene caption examples in Fig. A-10.
Figure 10: Example visualization of grounded scene caption dataset. The “[ ]” symbol marks each grounded phrase. The color of the phrase corresponds to the segmented point cloud.
Data generation prompts Fig. A-11 displays the object caption prompts, including the VLM captioning prompt [38] and the subsequent condensation of the caption into a descriptive phrase using ChatGPT. These object captions are then compiled to create local scene captions, as shown in Fig. A-12, while retaining their object IDs. Finally, spatial relationships derived from program-generated code are incorporated using the prompt depicted in Fig. A-13.
Figure 11: Object caption prompt for VLM and ChatGPT (used in Step-1 of data generation). The above is the prompt for CogVLM that utilizes the object labels from ScanNet-200. Below is the prompt for ChatGPT input, where “System” denotes the system prompt and “User” denotes user input.
Figure 12: GPT-4 prompt for scene caption (used in Step 2 of data generation). “System” denotes the system prompt and “User” denotes user input for GPT-4.
Figure 13: GPT-4 prompt of including rule-based relation (used in Step 3 of data generation). “System” denotes the system prompt and “User” denotes user input for GPT-4.
10 Task-specific Instruction-following Templates
Grounded 3D-LLM simultaneously undertakes vision tasks—such as 3D visual grounding, 3D object detection—and language tasks, like question answering. Each task demands specific, diverse templates to enable the LLM to naturally provide scene-aware answers tailored to particular dataset specifications. We show several instruction-following templates in the following figures.
Object detection. The detection tasks employ various question templates for each category, along with three answer templates to accommodate various detection outcomes – single, multiple, or no objects detected – as illustrated in Fig. A-14.
Figure 14: Instruction-following templates for object detection.
Single and multi-object grounding. The language grounding’s question template, as depicted in Fig. A-15, directs the model to identify referred objects. For the ScanRefer grounding dataset, the answer templates indicate single-object grounding. The answer templates for the Multi3DRef dataset accommodate various grounding scenarios—single, multiple, or no grounded objects.
Figure 15: Instruction-following templates for single- or multi- object grounding.
3D Question answering. For ScanQA, we append suffixes to brief outputs (Fig. A-16) to indicate the outputs with only phrases or word output as the dataset annotation guideline.
Figure 16: Instruction-following templates for ScanQA.
Figure 17: Instruction-following templates for Scan2Cap.
Dense captioning. Question templates for Scan2Cap require the model to describe object appearance followed by spatial relations, as illustrated in Fig. A-17.
11 Extension to embodied dialogue and embodied planning
Figure 18: GPT-4 Prompt of embodied dialogue. “System” denotes the system prompt and “User” denotes user input for GPT-4. “{Grounded scene caption}” is filled with our grounded scene caption data.
Figure 19: GPT-4 Prompt of embodied planning. “System” denotes the system prompt and “User” denotes user input for GPT-4. “{Grounded scene caption}” is filled with our grounded scene caption data.
Fig A-18 and A-19 list the prompts for grounded scene caption-based embodied dialogue and planning. The training dataset includes 11.9K dialogue examples and 4.4K planning examples.
Result visualization. As there is no well-defined benchmark for evaluating both phrase grounding and language understanding, we directly visualize the results of the jointly trained Grounded 3D-LLM for embodied dialogue (Fig. 20) and embodied planning (Fig. 21). These examples demonstrate the potential of Grounded 3D-LLM’s grounding capability to enhance robot navigation and manipulation.
Figure 20: Result visualization for embodied dialogue. Please zoom in for better visualization.
Figure 21: Result visualization for embodied planning. Please zoom in for better visualization.